# On the Learnability of Watermarks for Language Models

### v2

@@ Appendix B Full sampling-based watermark distillation results table
DecT WD-S DecT WD-S DecT DecS WD-S DecT DecS WD-S DecT DecS WD-S
KGW $\delta=2$ 4.4e-18 3.6e-06 1.000 0.995 16.4 32.8 51.9 0.04 0.04 0.02 0.89 0.81 0.65
$\delta=1$ 1.0e-04 2.0e-02 0.992 0.874 12.6 26.2 41.3 0.03 0.03 0.02 0.91 0.84 0.72
Aar $k=2$ 1.3e-75 2.9e-14 0.999 0.993 6.5 7.0 19.8 0.34 0.53 0.23 0.28 0.10 0.52
$k=3$ 4.8e-73 1.6e-02 1.000 0.866 9.5 15.1 29.0 0.14 0.27 0.10 0.74 0.46 0.68
$k=4$ 4.0e-72 3.0e-01 0.999 0.631 10.7 20.5 33.8 0.09 0.12 0.05 0.86 0.76 0.75
KTH $s=1$ 1.0e-04 (596) 1.0e-04 (456) 0.994 0.989 10.5 22.2 34.0 0.03 0.03 0.02 0.88 0.72 0.81
$s=2$ 1.0e-04 (600) 4.0e-04 (445) 0.998 0.972 10.7 21.9 26.3 0.03 0.03 0.02 0.91 0.79 0.84
$s=4$ 1.0e-04 (593) 2.2e-03 (436) 0.995 0.955 10.5 22.2 24.7 0.03 0.03 0.02 0.91 0.84 0.85
$s=256$ 1.0e-04 (600) 2.1e-03 (437) 0.996 0.955 10.6 22.2 25.7 0.03 0.03 0.02 0.92 0.85 0.88
Base student 5.0e-01 0.500 21.7 0.03 0.88
Table 4: Full version of Table 2, containing results for sampling-based watermark distillation experiments, including the seq-rep-3 metric.
Appendix C Llama 2 7B sampling-based watermark distillation experiments
@@
We run additional experiments on sampling-based watermark distillation using Llama 2 7B as both the teacher model and student model.
First, we use Llama 2 7B with a decoding-based watermarking strategy to generate 640k watermarked samples of length 256 tokens, prompted with 50-token prefixes from OpenWebText. Then, we fine-tune Llama 2 7B on the watermarked samples for 1 epoch of 5,000 steps, with a batch size of 128, sequence length of 256 tokens, maximal learning rate of 1e-5, and cosine learning rate decay with a linear warmup for the first 500 steps. Each training run took approximately 4 hours on 4 NVIDIA A100 80GB GPUs.
Results using the same evaluation procedure and metrics in §4.3 are shown in Table 5. The results exhibit similar trends as logits-based watermark distillation with Llama 2 7B and sampling-based watermark distillation with Llama 2 7B as the teacher model and Pythia 1.4B as the student model (§5), showing that sampling-based watermark distillation can also be used when the teacher and student models are the same.
Using Llama 2 7B as both the teacher and student model allows for a fairer comparison with logits-based watermark distillation (compare Tables 1 and 5). However, we note that logits-based and sampling-based watermark distillation have key differences, which means that they should not be compared solely based on efficacy to determine which method is better. Logits-based distillation requires white-box access to the teacher model and that the teacher and student models share the same tokenizer and vocabulary. It is also convenient if the teacher and student models share the same model architecture. On the other hand, sampling-based distillation only requires black-box access to samples from the teacher model, and the teacher and student models can have different tokenizers and vocabularies. Also, sampling-based distillation requires autoregressive sampling from the teacher model, which is slow, whereas logits-based distillation does not require sampling and is faster overall. So, logits-based and sampling-based distillation are each suitable and applicable for different settings, so neither is strictly better than the other in all scenarios.
@@
DecT WD-S DecT WD-S DecT WD-S DecT WD-S DecT WD-S
KGW $\delta=2$ 4.4e-18 8.5e-07 1.000 0.996 16.4 19.7 0.04 0.04 0.89 0.85
$\delta=1$ 1.0e-05 2.0e-02 0.992 0.869 12.5 15.4 0.03 0.03 0.91 0.89
Aar $k=2$ 1.3e-75 5.1e-17 0.999 0.984 6.5 7.6 0.34 0.34 0.28 0.29
$k=3$ 4.8e-73 8.5e-03 1.000 0.873 9.5 10.5 0.14 0.17 0.74 0.66
$k=4$ 4.0e-72 2.6e-01 0.999 0.651 10.7 11.9 0.09 0.10 0.86 0.83
KTH $s=1$ 1.0e-04 (596) 1.0e-04 (560) 0.994 0.998 10.5 15.3 0.03 0.03 0.88 0.86
$s=2$ 1.0e-04 (600) 1.0e-04 (526) 0.998 0.994 10.7 13.5 0.03 0.04 0.91 0.87
$s=4$ 1.0e-04 (593) 1.0e-04 (485) 0.995 0.984 10.5 12.6 0.03 0.04 0.91 0.87
$s=256$ 1.0e-04 (600) 1.0e-04 (455) 0.996 0.968 10.6 11.3 0.03 0.04 0.92 0.90
Base student 5.0e-01 0.500 11.7 0.03 0.92
@@
Table 5: Results of sampling-based watermark distillation experiments using Llama 2 7B as both the teacher and student model.
Appendix D Sample complexity of sampling-based watermark distillation
We run additional experiments to investigate the sample complexity of sampling-based watermark distillation. We run sampling-based watermark distillation on varying numbers of training samples and compute watermark detection p-values.
We use Llama 2 7B as the teacher model and Pythia 1.4B as the student model. Each training sample is 256 tokens long (using the Llama 2 tokenizer). We fine-tune Pythia 1.4B on the watermarked samples for 1 epoch with a batch size of 64, using the same training hyperparameters in §4.2. Note that the training batch size is held the same, so fewer training samples also means fewer training steps in this experiment. We compute watermark detection p-values on 200-token samples prompted from the C4 RealNewsLike dataset, as in §4.3.
Results are shown in Table 6 and Figure 5. The results show that as the number of training samples increases, the watermark is learned more strongly, as indicated by the smaller p-values. However, even at small numbers of training samples, the watermark is still learned, just to a lesser degree (the p-values are still smaller than 0.5). Overall, we note that sample efficiency is a continuous spectrum. More samples are helpful for learnability but not always necessarily crucial, and sample efficiency varies across watermarking strategies and hyperparameters.
number of training samples
Model 20k 40k 80k 160k 320k 640k
KGW $\delta=2$ 4.1e-02 1.4e-02 2.4e-03 5.2e-04 5.1e-05 3.6e-06
Aar $k=2$ 7.4e-02 9.4e-03 3.5e-04 1.0e-06 5.1e-10 2.9e-14
KTH $s=1$ 7.4e-02 4.2e-02 1.0e-02 1.1e-03 6.0e-04 1.0e-04
Table 6: p-values of generations from sampling-based distilled Pythia 1.4B models using varying numbers of training samples, each 256 tokens long. The training batch size is held the same, so fewer training samples means fewer training steps. We observe that as the number of training samples increases, the watermark is learned more strongly. The watermark is still learned at small numbers of training samples, just to a lesser degree.
Figure 5: p-values of generations from sampling-based distilled Pythia 1.4B models plotted against varying numbers of training samples, each 256 tokens long. As the number of training samples increases, the p-values become smaller.
Appendix E Mixing different keys for sampling-based watermark distillation
We run additional experiments to investigate the effect on sampling-based watermark distillation if the training samples are not all watermarked using the same key. Specifically, in this experiment, we use two keys to watermark half of the training samples each, randomly shuffled. As in Appendix D, we also vary the number of training samples. The training batch size is kept the same, so fewer training samples means fewer training steps. Using the same training details as in §4.2, we use Llama 2 7B as the teacher model and Pythia 1.4B as the student model.
Results are shown in Table 7 and Figure 6. The results show that using different keys to watermark the training samples hinders watermark learning, as indicated by the larger p-values when training on samples using two different keys. However, learnability is not completely prevented, as the p-values are still much smaller than 0.5 and get smaller as the number of training samples increases.
This suggests that using multiple watermark keys for generation may be a potential defense for language model providers against spoofing attacks. However, if multiple watermark keys are used for generation, then all of those keys will need to be tested at detection, so multiple testing correction will need to be performed to obtain accurate p-values. So, the p-values for watermarked text will be larger, and this defense will slightly reduce the statistical power and increase the false negative rate of watermark detection.
number of training samples
Model 20k 40k 80k 160k 320k 640k
KGW $\delta=2$ (one key) 4.1e-02 1.4e-02 2.4e-03 5.2e-04 5.1e-05 3.6e-06
KGW $\delta=2$ (two keys) 1.7e-01 1.4e-01 7.8e-02 5.7e-02 2.9e-02 1.4e-02
Aar $k=2$ (one key) 7.4e-02 9.4e-03 3.5e-04 1.0e-06 5.1e-10 2.9e-14
Aar $k=2$ (two keys) 4.2e-01 2.3e-01 1.2e-01 4.0e-02 1.0e-02 1.2e-03
Table 7: p-values of generations from sampling-based Pythia 1.4B models when the training samples are watermarked using only one key versus two different keys. We also vary the number of training samples, each 256 tokens long. We observe that using two different keys hinders watermark learning, as indicated by the larger p-values. However, learning is not completely prevented, as the p-values are still smaller than 0.5 and decrease as the number of training samples increases.
Figure 6: p-values of generations from sampling-based distilled Pythia 1.4B models against varying number of training samples. We compare watermarking the training samples using only one key versus two different keys. We observe that using two different keys hinders watermark learning, as indicated by the larger p-values. However, learning is not completely prevented, as the p-values are still smaller than 0.5 and decrease as the number of training samples increases.
Appendix F Additional datasets
We run evaluations on the additional datasets of Wikipedia articles (Foundation, 2022) and arXiv papers (Cohan et al., 2018). The evaluation procedure and metrics are the same as in §4.3, except for the dataset. We evaluate 5k 200-token completions from 50-token prompts.
Tables 8 and 9 show the results on Wikipedia articles and arXiv papers, respectively, for logits-based watermark distillation using Llama 2 7B. Tables 10 and 11 show the results on Wikipedia articles and arXiv papers, respectively, for sampling-based watermark distillation using Llama 2 7B as both the teacher and student model. Tables 12 and 13 show the results on Wikipedia articles and arXiv papers, respectively, for sampling-based watermark distillation using Llama 2 7B as the teacher model and Pythia 1.4B as the student model.
We observe similar trends on these additional datasets to the main evaluation on the C4 RealNewsLike dataset in §5.
@@ Appendix F Additional datasets
DecT WD-L DecT WD-L DecT WD-L DecT WD-L DecT WD-L
KGW $\delta=2$ 1.5e-17 1.9e-07 1.000 0.993 18.3 19.8 0.07 0.05 0.73 0.77
$\delta=1$ 2.7e-05 1.4e-02 0.985 0.893 13.5 13.9 0.05 0.05 0.78 0.83
Aar $k=2$ 1.6e-66 6.8e-07 1.000 0.975 6.2 11.2 0.36 0.11 0.27 0.77
$k=3$ 2.3e-64 2.0e-01 1.000 0.711 8.8 12.3 0.22 0.06 0.52 0.81
$k=4$ 7.2e-64 4.1e-01 1.000 0.562 10.1 13.0 0.14 0.05 0.66 0.81
KTH $s=1$ 1.0e-04 (564) 1.0e-04 (530) 0.996 0.996 10.4 16.6 0.06 0.06 0.79 0.68
$s=2$ 1.0e-04 (566) 1.0e-04 (466) 0.993 0.982 10.9 17.8 0.05 0.06 0.80 0.71
$s=4$ 1.0e-04 (567) 5.7e-03 (432) 0.995 0.923 10.4 14.4 0.05 0.06 0.82 0.71
$s=256$ 1.0e-04 (568) 1.1e-01 (420) 0.997 0.808 10.7 11.0 0.05 0.07 0.82 0.73
Base student 5.0e-01 0.500 12.0 0.05 0.82
Table 8: Results of Llama 2 7B logits-based watermark distillation experiments, evaluating on Wikipedia articles.
@@
DecT WD-L DecT WD-L DecT WD-L DecT WD-L DecT WD-L
KGW $\delta=2$ 2.9e-25 2.4e-07 1.000 0.994 41.0 40.6 0.07 0.04 0.43 0.50
$\delta=1$ 1.9e-07 1.4e-02 0.996 0.863 30.5 30.8 0.05 0.04 0.54 0.63
Aar $k=2$ 3.1e-106 5.4e-06 0.999 0.976 7.5 25.0 0.50 0.10 0.09 0.50
$k=3$ 1.1e-106 2.8e-01 0.999 0.657 15.4 28.6 0.30 0.05 0.26 0.54
$k=4$ 2.7e-108 4.6e-01 0.998 0.545 21.0 29.4 0.19 0.04 0.42 0.51
KTH $s=1$ 1.0e-04 (700) 1.0e-04 (643) 1.000 0.994 22.8 38.6 0.05 0.07 0.43 0.26
$s=2$ 1.0e-04 (711) 1.0e-04 (515) 1.000 0.986 24.5 37.8 0.05 0.08 0.48 0.29
$s=4$ 1.0e-04 (701) 1.0e-04 (453) 1.000 0.953 22.8 30.6 0.05 0.09 0.54 0.36
$s=256$ 1.0e-04 (709) 3.6e-02 (426) 1.000 0.878 23.2 20.1 0.05 0.10 0.56 0.52
Base student 5.0e-01 0.500 27.7 0.04 0.62
Table 9: Results of Llama 2 7B logits-based watermark distillation experiments, evaluating on arXiv papers.
[…]
DecT WD-S DecT WD-S DecT WD-S DecT WD-S DecT WD-S
KGW $\delta=2$ 1.5e-17 5.1e-05 1.000 0.982 18.3 18.6 0.07 0.06 0.73 0.80
$\delta=1$ 2.7e-05 4.6e-02 0.985 0.814 13.5 13.7 0.05 0.05 0.78 0.84
Aar $k=2$ 1.6e-66 2.8e-07 1.000 0.946 6.2 7.0 0.36 0.33 0.27 0.35
$k=3$ 2.3e-64 6.5e-02 1.000 0.780 8.8 9.3 0.22 0.24 0.52 0.46
$k=4$ 7.2e-64 3.3e-01 1.000 0.606 10.1 10.5 0.14 0.14 0.66 0.70
KTH $s=1$ 1.0e-04 (564) 1.0e-04 (519) 0.996 0.994 10.4 15.1 0.06 0.05 0.79 0.80
$s=2$ 1.0e-04 (566) 1.0e-04 (490) 0.993 0.984 10.9 13.6 0.05 0.05 0.80 0.82
$s=4$ 1.0e-04 (567) 1.0e-04 (459) 0.995 0.964 10.4 12.4 0.05 0.06 0.82 0.84
$s=256$ 1.0e-04 (568) 1.8e-03 (436) 0.997 0.924 10.7 11.1 0.05 0.06 0.82 0.84
Base student 5.0e-01 0.500 12.0 0.05 0.82
Table 10: Results of Llama 2 7B sampling-based watermark distillation experiments, evaluating on Wikipedia articles.
[…]
DecT WD-S DecT WD-S DecT WD-S DecT WD-S DecT WD-S
KGW $\delta=2$ 2.9e-25 4.2e-05 1.000 0.975 41.0 40.9 0.07 0.06 0.43 0.41
$\delta=1$ 1.9e-07 5.7e-02 0.996 0.743 30.5 31.4 0.05 0.05 0.54 0.45
Aar $k=2$ 3.1e-106 1.6e-06 0.999 0.883 7.5 8.5 0.50 0.51 0.09 0.06
$k=3$ 1.1e-106 1.1e-01 0.999 0.720 15.4 15.3 0.30 0.32 0.26 0.17
$k=4$ 2.7e-108 3.9e-01 0.998 0.577 21.0 22.4 0.19 0.20 0.42 0.35
KTH $s=1$ 1.0e-04 (700) 1.0e-04 (603) 1.000 0.999 22.8 30.0 0.05 0.05 0.43 0.38
$s=2$ 1.0e-04 (711) 1.0e-04 (557) 1.000 0.996 24.5 26.5 0.05 0.05 0.48 0.42
$s=4$ 1.0e-04 (701) 1.0e-04 (505) 1.000 0.993 22.8 24.4 0.05 0.06 0.54 0.42
$s=256$ 1.0e-04 (709) 1.0e-04 (462) 1.000 0.975 23.2 20.5 0.05 0.06 0.56 0.48
Base student 5.0e-01 0.500 27.7 0.04 0.62
Table 11: Results of Llama 2 7B sampling-based watermark distillation experiments, evaluating on arXiv papers.
[…]
DecT WD-S DecT WD-S DecT DecS WD-S DecT DecS WD-S DecT DecS WD-S
KGW $\delta=2$ 1.5e-17 2.7e-05 1.000 0.988 18.3 26.4 50.5 0.07 0.05 0.03 0.73 0.84 0.31
$\delta=1$ 2.7e-05 2.9e-02 0.985 0.854 13.5 21.2 40.9 0.05 0.04 0.02 0.78 0.87 0.34
Aar $k=2$ 1.6e-66 3.4e-09 1.000 0.974 6.2 7.3 18.4 0.36 0.46 0.24 0.27 0.20 0.30
$k=3$ 2.3e-64 4.2e-02 1.000 0.817 8.8 12.5 26.6 0.22 0.27 0.13 0.52 0.53 0.34
$k=4$ 7.2e-64 3.1e-01 1.000 0.627 10.1 16.6 33.1 0.14 0.14 0.06 0.66 0.78 0.36
KTH $s=1$ 1.0e-04 (564) 4.0e-04 (443) 0.996 0.975 10.4 18.3 30.5 0.06 0.04 0.03 0.79 0.84 0.74
$s=2$ 1.0e-04 (566) 5.7e-03 (432) 0.993 0.928 10.9 18.3 24.5 0.05 0.04 0.03 0.80 0.87 0.83
$s=4$ 1.0e-04 (567) 1.7e-02 (428) 0.995 0.911 10.4 18.4 23.4 0.05 0.04 0.04 0.82 0.90 0.81
$s=256$ 1.0e-04 (568) 1.5e-02 (428) 0.997 0.914 10.7 18.3 24.1 0.05 0.04 0.03 0.82 0.89 0.82
Base student 5.0e-01 0.500 20.2 0.04 0.89
@@
DecT WD-S DecT WD-S DecT DecS WD-S DecT DecS WD-S DecT DecS WD-S
KGW $\delta=2$ 2.9e-25 5.1e-05 1.000 0.976 41.0 41.7 68.4 0.07 0.06 0.03 0.43 0.52 0.07
$\delta=1$ 1.9e-07 4.1e-02 0.996 0.783 30.5 33.8 57.1 0.05 0.04 0.02 0.54 0.54 0.08
Aar $k=2$ 3.1e-106 1.5e-09 0.999 0.954 7.5 7.6 18.4 0.50 0.54 0.32 0.09 0.09 0.06
$k=3$ 1.1e-106 5.5e-02 0.999 0.796 15.4 17.7 33.9 0.30 0.29 0.16 0.26 0.31 0.07
$k=4$ 2.7e-108 3.3e-01 0.998 0.621 21.0 24.4 43.3 0.19 0.17 0.09 0.42 0.43 0.08
KTH $s=1$ 1.0e-04 (700) 1.0e-04 (451) 1.000 0.979 22.8 29.7 39.7 0.05 0.04 0.03 0.43 0.40 0.37
$s=2$ 1.0e-04 (711) 5.0e-04 (444) 1.000 0.972 24.5 28.2 34.9 0.05 0.04 0.04 0.48 0.43 0.37
$s=4$ 1.0e-04 (701) 2.8e-03 (437) 1.000 0.956 22.8 29.0 32.6 0.05 0.04 0.04 0.54 0.48 0.36
$s=256$ 1.0e-04 (709) 2.9e-03 (437) 1.000 0.954 23.2 28.3 33.1 0.05 0.04 0.04 0.56 0.49 0.36
Base student 5.0e-01 0.500 31.9 0.04 0.57

### v3

@@ Appendix C Pythia sampling-based watermark distillation experiments
Watermark
Decoding (T)
Sampling
Decoding (T)
Sampling
Decoding (T)
Decoding (S)
Sampling
Decoding (T)
Decoding (S)
Sampling
KGW $k=0,\delta=2$ 6e-16 2e-17 1.00 1.00 17.5 31.1 58.8 0.05 0.07 0.03
$k=1,\delta=2$ 4e-18 4e-06 1.00 0.99 16.5 34.7 56.0 0.04 0.04 0.02
$k=2,\delta=2$ 9e-18 1e-01 1.00 0.75 16.8 36.9 58.7 0.03 0.03 0.01
$k=0,\delta=1$ 5e-04 3e-04 0.98 0.99 13.0 28.2 47.9 0.03 0.03 0.02
$k=1,\delta=1$ 1e-05 2e-02 0.99 0.87 12.7 27.7 44.6 0.03 0.03 0.02
Aar $k=2$ 1e-75 3e-14 1.00 0.99 6.5 7.2 20.6 0.34 0.53 0.23
$k=3$ 5e-73 1e-02 1.00 0.88 9.5 15.8 31.2 0.14 0.27 0.10
$k=4$ 4e-72 3e-01 1.00 0.63 10.7 21.6 37.0 0.09 0.12 0.05
KTH $s=1$ 1e-04 (-593) 1e-04 (-457) 1.00 0.99 10.5 23.4 35.9 0.03 0.03 0.02
$s=2$ 1e-04 (-596) 4e-04 (-445) 1.00 0.97 10.7 23.1 28.2 0.03 0.03 0.02
$s=4$ 1e-04 (-594) 2e-03 (-437) 1.00 0.96 10.6 23.0 26.2 0.03 0.03 0.02
$s=256$ 1e-04 (-594) 2e-03 (-436) 1.00 0.95 10.8 23.4 27.6 0.03 0.03 0.02
Base student 5e-01 0.50 26.4 0.03
Table 3: Results for the Pythia 1.4B sampling-based watermark distillation experiments. Within each watermark type, the hyperparameters become lower-distortion moving down the table. Higher-distortion watermarks are successfully learned with small p-values and high detectability. Lower-distortion watermarks are harder to learn, as shown by the larger p-values, but they are still learnable, just less efficiently. The results indicate that sampling-based watermark distillation is also effective when the teacher and student models have different tokenizers and sizes.
Appendix D Sample complexity of watermark distillation
In this experiment, we investigate the sample complexity of logit-based and sampling-based watermark distillation.
Experimental setup. We use one hyperparameter setting from each watermark type: KGW $k=1,\gamma=0.25,\delta=2$, Aar $k=2$, and KTH $s=4$. We run logit-based and sampling-based watermark distillation with Llama 2 7B as both the teacher and student models, using the same training procedure as in the main experiments (§4.2), except with varying numbers of tokens trained on. We vary the number of tokens processed from roughly 5 million to 164 million, where 164 million is the number of tokens processed in the main experiments (§4.2). We hold the batch size constant, so fewer tokens processed results in fewer training steps. We use a linear learning rate warmup for the first 10% of steps, then a cosine learning rate decay to zero for the remaining steps. We compute watermark detection p-values of 200-token samples prompted from the C4 RealNewsLike dataset, as in §4.3.
Results. Results are shown in Figure 5. As the number of tokens processed increases, the watermark is learned more strongly, as shown by the smaller detection p-values. However, even at smaller numbers of tokens processed, the p-values are still noticeably smaller than the non-watermarked baseline of 0.5, showing that the watermark is still learned, just to a lesser degree. Overall, sample complexity is a continuous spectrum. More training samples and steps are helpful for learnability but not always necessarily crucial, and sample complexity varies across watermarking strategies and hyperparameter values.
(a) Logit-based distillation sample complexity
(b) Sampling-based distillation sample complexity
Figure 5: Watermark detection p-values of generations from logit-based (a) and sampling-based (b) distilled Llama 2 7B models trained on varying numbers of tokens. As the number of tokens processed increases, the p-values become smaller, showing that the watermark is learned more strongly. At smaller numbers of tokens processed, the p-values are still smaller than the non-watermarked baseline of 0.5, indicating that the watermark is still learned, albeit less strongly.
Appendix E Mixing two keys for sampling-based watermark distillation
In this experiment, we investigate the effect on sampling-based watermark distillation if the training samples are not all watermarked using the same key.
Experimental setup. We use one hyperparameter setting from each watermark type: KGW $k=1,\gamma=0.25,\delta=2$, Aar $k=2$, and KTH $s=1$. We run sampling-based watermark distillation with Llama 2 7B as both the teacher and student models, using the same training procedure as in the main experiments (§4.2). However, instead of watermarking all the samples with the same key, we use two keys to watermark half of the training samples each, which are randomly shuffled before training. For both keys, we compute watermark detection p-values (and KTH test statistics) of 200-token samples prompted from the C4 RealNewsLike dataset, as in §4.3. We compare against the p-values achieved by sampling-based distillation when all samples are watermarked using the same key (from Table 1).
Results. Results are shown in Table 4. Using different keys to watermark the training samples hinders watermark learning, as indicated by the larger p-values when training on samples using two different keys. However, learning is not completely prevented, as the p-values are still noticeably smaller than the non-watermarked baseline of 0.5. Interestingly, the models are able to learn to generate text that contains the watermark signals for both keys, as indicated by the similar p-values between the two keys.
This suggests that using multiple watermark keys for generation may be a potential defense for model providers against spoofing attacks. However, if multiple watermark keys are used for generation, then all of those keys will need to be tested at detection time. Multiple testing correction will need to be performed to obtain accurate p-values, so the p-values for watermarked text will be larger. So, this defense will slightly reduce the statistical power and increase the false negative rate of watermark detection.
Different keys
Model Same key Key 1 Key 2
KGW $k=1,\delta=2$ 8e-07 1e-02 3e-02
Aar $k=2$ 3e-17 3e-03 4e-03
KTH $s=1$ 1e-04 (-561) 8e-02 (-423) 1e-04 (-489)
Table 4: Watermark detection p-values (and KTH test statistics) of generations from sampling-based distilled Llama 2 7B models when the training samples are all watermarked using the same key versus two different keys. Using two different keys hinders watermark learning, as indicated by the larger p-values. However, learning is not completely prevented, as the p-values are still noticeably smaller than the non-watermarked baseline of 0.5.
Appendix F Additional datasets
We run evaluations on the additional datasets of Wikipedia articles (Foundation, 2022) and arXiv papers (Cohan et al., 2018). The evaluation procedure and metrics are the same as in §4.3, except for the dataset. We evaluate 5,000 200-token completions from 50-token prompts.
@@ Appendix F Additional datasets
Tables 5 and 6 show the results on Wikipedia articles and arXiv papers, respectively, for logit-based and sampling-based watermark distillation using Llama 2 7B as both the teacher and student models (§4.2). Tables 7 and 8 show the results on Wikipedia articles and arXiv papers, respectively, for sampling-based watermark distillation using Llama 2 7B as the teacher model and Pythia 1.4B as the student model (Appendix C).
Results on Wikipedia articles and arXiv papers exhibit similar trends as the main evaluation on the C4 RealNewsLike dataset (Tables 1 and 3), indicating that watermark distillation is relatively robust to domain shifts.
@@
Watermark
Decoding
Logit
Sampling
Decoding
Logit
Sampling
Decoding
Logit
Sampling
Decoding
Logit
Sampling
KGW $k=0,\delta=2$ 5e-17 3e-19 2e-15 1.00 1.00 1.00 19.1 17.4 20.0 0.08 0.09 0.07
$k=1,\delta=2$ 2e-17 1e-07 3e-05 1.00 0.99 0.98 18.2 19.4 19.4 0.07 0.05 0.06
$k=2,\delta=2$ 3e-17 1e-01 2e-01 1.00 0.67 0.64 19.3 21.3 19.3 0.05 0.04 0.04
$k=0,\delta=1$ 5e-04 1e-05 1e-03 0.96 0.98 0.95 14.1 13.6 14.9 0.05 0.06 0.05
$k=1,\delta=1$ 3e-05 1e-02 5e-02 0.98 0.89 0.82 12.9 13.8 14.0 0.06 0.05 0.05
Aar $k=2$ 2e-66 5e-07 2e-07 1.00 0.98 0.95 6.2 11.4 7.0 0.36 0.11 0.33
$k=3$ 2e-64 2e-01 7e-02 1.00 0.71 0.78 8.8 12.2 9.0 0.22 0.06 0.24
$k=4$ 7e-64 4e-01 3e-01 1.00 0.55 0.60 10.0 12.7 10.3 0.14 0.05 0.14
KTH $s=1$ 1e-04 (-570) 1e-04 (-538) 1e-04 (-525) 1.00 0.99 1.00 10.3 16.4 14.9 0.06 0.06 0.05
$s=2$ 1e-04 (-574) 1e-04 (-468) 1e-04 (-494) 1.00 0.98 0.98 10.7 17.9 13.6 0.06 0.06 0.05
$s=4$ 1e-04 (-574) 5e-03 (-433) 1e-04 (-463) 1.00 0.93 0.97 10.6 14.3 12.5 0.05 0.06 0.05
$s=256$ 1e-04 (-573) 9e-02 (-421) 2e-03 (-437) 1.00 0.82 0.93 10.7 11.1 11.0 0.05 0.07 0.05
Base student 5e-01 0.50 12.0 0.05
@@
Table 5: Results of Llama 2 7B logit-based and sampling-based watermark distillation experiments, evaluating on Wikipedia articles.
@@
Watermark
Decoding
Logit
Sampling
Decoding
Logit
Sampling
Decoding
Logit
Sampling
Decoding
Logit
Sampling
KGW $k=0,\delta=2$ 6e-33 2e-31 1e-28 1.00 1.00 1.00 32.1 31.6 35.6 0.13 0.14 0.11
$k=1,\delta=2$ 5e-25 3e-07 5e-05 1.00 0.99 0.98 40.4 40.0 40.9 0.07 0.04 0.06
$k=2,\delta=2$ 3e-24 2e-01 3e-01 1.00 0.65 0.63 42.6 49.0 46.2 0.05 0.03 0.04
$k=0,\delta=1$ 3e-08 7e-09 8e-07 0.99 1.00 0.99 30.1 29.9 32.4 0.05 0.06 0.05
$k=1,\delta=1$ 2e-07 1e-02 6e-02 1.00 0.86 0.74 29.8 30.1 32.4 0.05 0.04 0.05
Aar $k=2$ 3e-106 5e-06 6e-07 1.00 0.97 0.88 7.5 25.3 8.4 0.50 0.09 0.51
$k=3$ 1e-106 3e-01 1e-01 1.00 0.65 0.71 15.4 28.1 14.9 0.30 0.05 0.32
$k=4$ 3e-108 5e-01 4e-01 1.00 0.54 0.58 21.0 29.2 22.4 0.19 0.05 0.19
KTH $s=1$ 1e-04 (-703) 1e-04 (-641) 1e-04 (-601) 1.00 1.00 1.00 22.9 38.5 28.9 0.05 0.07 0.05
$s=2$ 1e-04 (-708) 1e-04 (-516) 1e-04 (-558) 1.00 0.98 1.00 23.8 37.3 26.2 0.05 0.08 0.05
$s=4$ 1e-04 (-700) 1e-04 (-454) 1e-04 (-503) 1.00 0.96 0.99 22.7 30.7 24.1 0.05 0.09 0.06
$s=256$ 1e-04 (-707) 3e-02 (-426) 1e-04 (-463) 1.00 0.86 0.98 23.6 20.1 20.6 0.04 0.10 0.06
Base student 5e-01 0.50 26.8 0.04
Table 6: Results of Llama 2 7B logit-based and sampling-based watermark distillation experiments, evaluating on arXiv papers.
@@
Watermark
Decoding (T)
Sampling
Decoding (T)
Sampling
Decoding (T)
Decoding (S)
Sampling
Decoding (T)
Decoding (S)
Sampling
KGW $k=0,\delta=2$ 5e-17 2e-17 1.00 1.00 19.1 26.7 59.3 0.08 0.08 0.04
$k=1,\delta=2$ 2e-17 3e-05 1.00 0.99 18.2 29.1 55.4 0.07 0.05 0.03
$k=2,\delta=2$ 3e-17 1e-01 1.00 0.68 19.3 30.5 58.9 0.05 0.04 0.02
$k=0,\delta=1$ 5e-04 3e-04 0.96 0.98 14.1 23.0 47.3 0.05 0.04 0.02
$k=1,\delta=1$ 3e-05 3e-02 0.98 0.85 12.9 23.0 44.4 0.06 0.04 0.02
Aar $k=2$ 2e-66 3e-09 1.00 0.97 6.2 7.7 19.5 0.36 0.46 0.24
$k=3$ 2e-64 4e-02 1.00 0.82 8.8 13.3 28.7 0.22 0.27 0.14
$k=4$ 7e-64 3e-01 1.00 0.62 10.0 17.7 36.0 0.14 0.14 0.06
KTH $s=1$ 1e-04 (-570) 4e-04 (-444) 1.00 0.97 10.3 19.5 33.1 0.06 0.04 0.03
$s=2$ 1e-04 (-574) 6e-03 (-433) 1.00 0.94 10.7 19.6 26.4 0.06 0.04 0.03
$s=4$ 1e-04 (-574) 1e-02 (-429) 1.00 0.92 10.6 19.7 25.5 0.05 0.04 0.03
$s=256$ 1e-04 (-573) 1e-02 (-429) 1.00 0.92 10.7 19.5 26.1 0.05 0.04 0.03
Base student 5e-01 0.50 21.1 0.04
@@
Watermark
Decoding (T)
Sampling
Decoding (T)
Sampling
Decoding (T)
Decoding (S)
Sampling
Decoding (T)
Decoding (S)
Sampling
KGW $k=0,\delta=2$ 6e-33 6e-21 1.00 1.00 32.1 34.4 72.3 0.13 0.10 0.05
$k=1,\delta=2$ 5e-25 5e-05 1.00 0.98 40.4 44.4 71.5 0.07 0.06 0.03
$k=2,\delta=2$ 3e-24 2e-01 1.00 0.70 42.6 49.2 78.3 0.05 0.04 0.02
$k=0,\delta=1$ 3e-08 9e-05 0.99 0.98 30.1 35.0 61.0 0.05 0.04 0.02
$k=1,\delta=1$ 2e-07 4e-02 1.00 0.78 29.8 36.1 57.8 0.05 0.04 0.02
Aar $k=2$ 3e-106 1e-09 1.00 0.96 7.5 7.7 19.3 0.50 0.54 0.32
$k=3$ 1e-106 5e-02 1.00 0.80 15.4 18.2 34.3 0.30 0.29 0.16
$k=4$ 3e-108 3e-01 1.00 0.63 21.0 25.2 44.6 0.19 0.17 0.08
KTH $s=1$ 1e-04 (-703) 1e-04 (-451) 1.00 0.98 22.9 30.7 41.7 0.05 0.04 0.03
$s=2$ 1e-04 (-708) 5e-04 (-445) 1.00 0.97 23.8 29.0 36.6 0.05 0.04 0.04
$s=4$ 1e-04 (-700) 3e-03 (-436) 1.00 0.96 22.7 29.3 33.6 0.05 0.04 0.04
$s=256$ 1e-04 (-707) 3e-03 (-437) 1.00 0.96 23.6 29.3 34.6 0.04 0.04 0.04
Base student 5e-01 0.50 32.8 0.04